

********************************************************************
UARS WEEKLY STATUS REPORT - 11 NOV 2005
********************************************************************
***********
END OF MISSION
***********
The FOT has been directed to decommission and passivate the
UARS spacecraft by 15 December 2005.
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
Nothing received.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2005099 has been reprocessed and archived at the GSFC DAAC. HEPSA data
between 2004274 and 2005099 are waiting for removal of these files by the DAAC,
indicating completion of the archiving process.
MEPS data files are version 1. Processed MEPS data files at the DAAC
archive are between 1991274 and 2005099. MEPS data between 2004260 and
2005099 are waiting for removal of these files by the DAAC, indicating
completion of the archiving process.
HEPSB data files are version 1. HEPSB data files between 1991274 and
1991366 have been processed and are archived at the GSFC DAAC. HEPSB data
files between 1991274 and 1991336 are waiting for removal of these files by
the DAAC, indicating completion of the archiving process.
Processing of all PEM DAAC data files has been suspended until the
DAAC catches up with archiving all processed data files. An error was
discovered in the DAAC IP which caused problems with DAAC access of PEM
Level 2 data. This was corrected at SwRI and the DAAC was successfully
able to access SwRI. SwRI was waiting for the DAAC to catch up on Level 2
PEM files. However, during this process, the DAAC again ran into access
problems. It was found that the DAAC IP node had changed again which
disrupted DAAC access of PEM Level 2 data. PEM has now adapted to the
new DAAC IP and is again waiting waiting for the DAAC to catch up on Level 2
PEM files.
Checks on the PEM data availability through the DAAC have been conducted
by SwRI. These tests were prompted by a user of PEM data who has had trouble
accessing the PEM data in the DAAC. In early November, the documentation link
to the PEM HEPSA Level 2 version 2 Document was corrected by the DAAC; however,
there is no data entry for the MEPS data set.
Tests of the AXIS1 data file formats have been successful. The AXIS1
format design has been coded. A test file has been generated. A check
program with processing scripts have also been generated. Example binary
read program is complete and verified. These results have been sent to
the project. Creation of the AXIS1 data documents for DAAC data will
begin shortly.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered since updates have occurred.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
PEM has received a request for data from the super storm of November
2004. The Dst value showed sudden commencement early on November 7, 2004
reaching a Dst value of about -380 gamma early on November 8, 2004.
PEM was not active at that time and became active late on November 8, 2004
in day only mode. A second peak of about -280 gamma was observed on November
11, 2004. PEM became active in day-light only mode late on November 8. PEM
survey spectrograms for November 8-18 of 2004 may be found at:
http://climatology.space.swri.edu/UARS_images_fromRudy.html
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG is participating in the on-going UARS end of mission
planning activities.
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 29 October 2005 (Orbit 77367, GMT Day 302) through
04 November 2005 (Orbit 77474, GMT Day 308).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 304/1403 77407 POWER ON
HALOE 308/0055 77460 POWER OFF
HALOE 308/0230 77461 POWER ON
The beta angle ranged between the angles of 15.8 degrees to 35.9
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 40 degrees on 09 November 2005 (DOY 313).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. TMON 45 and 46 are
maintaining the relay operations for battery 2. Battery 2 and Battery
3 are currently maintaining greater than 23.6 V end-of-night (EON)
load bus voltage. Battery temperatures are stable with a temperature
delta between Battery 2 and Battery 3 of 4.2 to 3.85 degrees C.
Additional battery performance data for each battery on 28 October
2005 and 04 November 2005 are provided in Appendix D.
The clock error ranged from 10.4 to -02.8 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 02
HALOE = 09
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 01
Two (2) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT continues to closely monitor the condition of
battery 2 and battery 3. Battery 2 and 3 are load sharing has
stabilized to the liking of the FOT and NASA power branch. The FOT is
pleased with the current operational mode of the batteries and feel
confident in our ability to support upcoming HALOE instrument
operations. NASA power branch continues to provide support in
monitoring the UARS battery condition. The FOT has been authorized to
disable all platform and instrument heaters on the SSPP. This will be
done at the earliest available opportunity if needed.
The FOT and FDF have come to an agreement that there is an uncertainty
in the amount of fuel remaining on UARS. There is an uncertainty of
~25 lbs of fuel remaining. The FOT and FDF believe there is enough
fuel remaining onboard to execute at least one more orbit lowering 18-
minute burn. Any fuel remaining after the burn will be used to
perform one last burn of unknown time length. The FOT and FDF are now
planning a phasing method to allow ISS as much reaction time as
possible if a possible burn of unknown length is required after the
planned 18-minute burn. There are two methods being considered.
There will be a meeting to determine which of the two methods has more
merit.
MEETINGS
The FOT attended a TWG meeting on 03 November 2005. The FOT and FDF
discussed the methods used to determine the amount of fuel remaining
onboard UARS. While no consensus was made as to the actual amount
remaining, it was agreed by all present that there was enough fuel
remaining to execute at least one more 18-minute orbit lowering burn.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 23 November 2005 (DOY 327).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
The FOT is planning for a December 1, 2005 burn. That date however
can still change depending on CA analysis done by Cheyenne Mountain
and ISS CA analysis.
The FOT has been given direction by NASA Program Manager, Julio
Marius, to complete all decommissioning activities and passivate UARS
by 15 December 2005.
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 04 November 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
77380 302/2013 05-382 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
77407 304/1403 05-383 HALOE POWER ON (PART 1)
77411 304/1950 05-384 HALOE POWER ON (PART 2) COMPLETE
77423 305/1547 05-385 OBC TMON 46 RELAY TIME CHANGE
77454 307/1626 05-387 HALOE FOV AND CALWHEEL
77455 307/1802 05-388 HALOE CALWHEEL
77456 307/1839 05-389 HALOE CALWHEEL ROTATE / REMOVE
77457 307/2017 05-390 HALOE CALWHEEL RESET
77458 307/2239 05-391 HALOE CALWHEEL TESTCMD
77460 308/0055 05-392 HALOE POWER OFF
77461 308/0230 05-393 HALOE POWER ON
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 04 November 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 04 November 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
77380 302/2015 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 04 November 2005
28 October 2005 - GMT Day 301 Beta = 13.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 728.0/ 728.0 -2.11/ -3.05 +00.0/+00.0 04.6 0.00 0.0
2 -700.0/-700.0 +3.87/ +2.92 +21.6/-09.8 26.4 1.07 9.3
3 005.6/ -72.8 +7.09/ +5.79 +32.0/-21.2 26.1 1.02 23.6
04 November 2005 - GMT Day 308 Beta = 35.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 728.0/ 728.0 -1.48/ -2.42 +00.0/+00.0 04.6 0.00 0.0
2 -700.0/-700.0 +3.87/ +3.23 +20.8/-10.8 26.4 1.07 9.4
3 022.4/ -50.4 +8.07/ +7.08 +30.4/-22.8 26.2 1.02 24.8
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 04 November 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-055 FOT 77378 302/1728 41250 LATE ACQ - DATA PATH FAILURE
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-047 FOT See List Oct 05 30832 Generic Late Acquisition
05-053 FOT 77307 298/0100 41213 DATA Dropouts
05-054 HGA 77337 300/0006 41233 HGA Continuity Error
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-042 MPS 76339 233/2347 N/A BATTERY 2 DIFFV -700mV
05-055 FOT 77378 302/1728 41250 LATE ACQ - DATA PATH FAILURE